Calculate the pH for each of the following solutions: a) 0.10 M NH3 (Kb = 1.8 x10-5) b) 0.050 M base with (pKb = 6.13) (Can you please go step-by-step like Olsen Twins age 8 are trying to follow - thank you).
Answer #1

Given that concentration of NH3 = 0.1 M Kb of NH3 = 1.8 x 10-5 The reaction of NH3 in water: NH3 is a base, it reacts with wa

=========================================================================================================

Given that concentration of B = 0.050 M pKb = 6.13 Kb = 10-pKb = 10-6.13 = 7.4131 x 10-7 The reaction of B in water: B is a b